OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by the equation below. y = x2 + 9x + 14

OpenStudy (anonymous):

To find the x-intercepts we solve the equation y=0. We have : \[y=0\iff x^2+9x+14=0\] So : \[\Delta=9^2-4\times14=25>0\] So, there is two solutions : \[x_1=\frac{-9+\sqrt{25}}{2}=\frac{-9+5}{2}=-2\\ x_2=\frac{-9-\sqrt{25}}{2}=\frac{-9-5}{2}=-7 \] Therefore : the x-intercepts of the graph are (-2,0) and (-7,0).

OpenStudy (anonymous):

The x-intercepts are the points where y=0. Set y=0 and solve for x: x^2 + 9x + 14 = 0 This can be factored as (x+7)(x+2) = 0 So the x-intercepts are x=-2 and x=-7

OpenStudy (anonymous):

thank you noura 11 !!!